# Yamaji effect in models of underdoped cuprates

### v1

@@ Abstract
Recent angle-dependent magnetoresistance measurements in underdoped cuprates have revealed compelling evidence for small hole pockets in the pseudogap regime, including observation of the Yamaji effect in HgBa2CuO4+δ (Chan et al., Nature Physics 10.1038/s41567-025-03032-2 (2025)). A key distinction between theories is their predicted Fermi volumes, measured as fractions of the square lattice Brillouin zone: $p/4$ per pocket for spin density wave (SDW) versus $p/8$ for fractionalized Fermi liquid (FL*), where $p$ is the hole doping. We calculate the $c$-axis magnetoresistance $\rho_{zz}(\theta,\phi)$ within the semiclassical Boltzmann formalism for both states, and using the ancilla layer model (ALM) for FL* in a single-band Hamiltonian. While the SDW Yamaji azimuthal angle $\theta$ is marginally smaller than the FL* Yamaji $\theta$, the most significant difference is the presence of a second Yamaji peak in the SDW theory near in-plane angle $\phi=45^{\circ}$; there is no such second peak in the observations in HgBa2CuO4+δ. Furthermore, we examine the role of anisotropic interlayer hopping present in LSCO, and predict the potential emergence of the Yamaji effect in LSCO at higher magnetic fields. Our results support the FL* interpretation of Fermi arcs in the pseudogap phase, and establish Yamaji angle measurements as a discriminatory tool between theoretical models.
@@ I Introduction
As the parent state from which superconductivity emerges, the pseudogap phase plays a central role in high-$T_{c}$ cuprates. Despite decades of intensive research, its microscopic origin remains one of the most prominent open questions in the field [1, 2, 3]. A key puzzle concerns the topology of the Fermi surface, which in the pseudogap regime manifests as an open “Fermi arc” rather than a closed, large Fermi surface. Whether these Fermi arcs represent the surviving segments of a large Fermi surface or instead form small Fermi pockets remains highly debated.
In principle, quantum oscillation measurements could discriminate between these scenarios by revealing the Fermi surface geometry directly. However, the very high upper critical fields near optimal doping make such measurements technically challenging. Recent progress in high-field transport measurements, particularly $c$-axis angle-dependent magnetoresistance (ADMR) [4, 5], provides a new opportunity to probe the Fermi pockets directly in the pseudogap phase without the presence of field-induced charge density wave order. Chan et al. [5] found that the Yamaji peak positions are consistent with small hole pockets of area $A_{\mathrm{FS}}=p/8$ (measured in units of the Brillouin zone area), rather than a large hole-like Fermi surface of area $A_{\mathrm{Large\,FS}}=(1+p)/2$, where $p$ is the hole doping level. In Ref. [5], the Fermi pocket is approximated as a simple ellipse with area $A_{\mathrm{pocket}}=p/8$, ignoring the Fermi-arc nature of the quasiparticle spectrum. As the spectral weight vanishes on the “back side” of the pocket, it is crucial to revisit whether the Yamaji peak can be reproduced within a more realistic microscopic model of the pseudogap state.
@@
One widely discussed interpretation attributes the Fermi arcs to a spin-density-wave (SDW) order with wavevector near $(\pi,\pi)$ [6, 7, 8, 9, 10, 11, 12]. In this picture, the SDW order coexists with a conventional Fermi liquid (FL) state possessing a large Fermi surface. The SDW order reconstructs the large Fermi surface by folding it with wavevector $(\pi,\pi)$, leading to small hole pockets. This framework provides a satisfactory description in extremely underdoped regime where robust long-range SDW order is present [13]. However, its applicability at higher doping levels is less clear, particularly in the finite-temperature regime where SDW order becomes short-ranged and strongly fluctuating [11].
An alternative scenario does not invoke static SDW order. Instead, it assumes that small Fermi surfaces coexist with a spin-liquid-like background, forming a fractionalized Fermi liquid (FL*) phase [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24]. In this case, the small Fermi pocket arises from the binding of holons and spinons into gauge-neutral electron-like quasiparticles.
Another popular approach to the pseudogap is the ‘phase fluctuation’ theory [25, 26, 27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 37, 38]. While this can describe the photoemission Fermi arc spectra in a limited range of temperatures above the superconducting $T_{c}$, it does not have a simple connection to the hole pockets which are the focus of our attention. Thus we focus on the SDW and FL* phases which are more consistent with the observation of the Yamaji effect.
@@
To check if ADMR can distinguish between the SDW and FL* scenarios, we calculate the $c$-axis magnetoresistance $\rho_{zz}(\theta,\phi)$ within the semiclassical Boltzmann formalism for both cases. A mean-field theory for the FL* state in a single band model, amenable to transport computations, is provided by the ancilla layer model (ALM) [21]. We show that the FL* framework reproduces the observed Yamaji peak as a function of the azimuthal angle $\theta$, at $p=0.1$ in HgBa2CuO4+δ [5]. For in-plane angle $\phi$ close to $0^{\circ}$, the value of the Yamaji $\theta$ is marginally smaller for the SDW theory, but the difference can be magnified by changes in the shape of the FL* hole pocket and the precise definition of the physical electron operator (see Appendix A). Moreover, there is a significant distinction between the two states at $\phi=45^{\circ}$. Here the SDW scenario predicts an additional peak at $\theta\approx 70^{\circ}$, which is absent in both the FL* calculation and in the current experiment [5]. The current experimental result at $\phi=45^{\circ}$ is closer to the FL* scenario than the SDW scenario. A more definitive conclusion can be reached at larger $\omega_{c}\tau$.
Our results support the FL* interpretation of Fermi arcs in the pseudogap phase, and demonstrate that Yamaji-angle measurements provide a powerful means of distinguishing between competing theoretical scenarios of the pseudogap. We propose future experiments to better distinguish the FL* and SDW scenario based on the data along the in-plane angle $\phi=45^{\circ}$ using higher magnetic field or cleaner sample.
Figure 1: (a) and (b) shows the Fermi pocket by (a) SDW and (b) FL* by ALM at doping $p=0.1$. In (b), the red color represents the contribution of the physical electron to the spectral weight on the Fermi surface, and purple color represents the contribution from the ancilla fermion $\psi_{1}$. The dark dashed line in (a) indicate the folded Brillion zone corresponding by SDW $(\pi,\pi)$. (c) An illustration of the ancilla layer model. In the one-band model considered in this work, $c$ and $\psi_{1}$ are the active electrically charged degrees of freedom, while the neutral $\psi_{2}$ fermions decouple and do not contribute to charge transport. (d) Schematic of the polar $\theta$ and azimuthal $\phi$ angles of the magnetic field $\bf B$. And an illustration of the caliper momentum $k_{\mathrm{cal}}$ along the in-plange magnetic field direction $\bf B_{\parallel}$.
@@
SDW metal: In this paper, we mainly consider the $Q=(\pi,\pi)$ SDW order, described by the mean field Hamiltonian:
$H_{\mathrm{SDW}}=\sum_{\mathbf{k},\sigma}\epsilon_{c}(\mathbf{k})c^{\dagger}_{\mathbf{k};\sigma}c_{\mathbf{k};\sigma}+\sum_{\mathbf{k},\sigma}\Delta c_{\mathbf{k};\sigma}^{\dagger}c_{\mathbf{k}+\mathbf{Q};\sigma}+\mathrm{h.c.}.$ (1)
where $\epsilon_{c}(\mathbf{k})$ denotes the dispersion of free fermions , $\mathbf{Q}=(\pi,\pi)$ is the SDW ordering wavevector, and $\Delta$ is the SDW order parameter. When $\Delta$ is finite, portions of the original Fermi surface connected by $Q$ are folded together, and a gap opens due to SDW ordering, leading to the formation of small closed pockets. We assume that the SDW order is strong enough to fully gap the electron-like pockets near $(\pi,0)$. The typical Fermi surfaces are shown in Fig. 1(a). In the folded smaller Brillouin zone (BZ), there are two independent Fermi pockets, each with an area of $A=p/4$ in units of the original BZ area, where the factor of 4 in the denominator accounts for the two pockets and two spin degeneracy.
FL* phase: The pseudogap phenomenon is robust across a range of temperatures and doping levels, suggesting that the gap opening and Fermi surface reconstruction may be independent of the symmetry breaking orders. Here we consider another picture that the doped holes (with density $p$) form small pockets on top of the spin moments (with density $1$) from the parent Mott insulator. The exact nature of the spin state for the local moments does not matter much for the hole pockets as long as there is no spin-rotation symmetry breaking. Therefore, we assume the spin moments form a symmetric spin liquid phase. The resulting metallic state is referred fractional Fermi liquid (FL*). The description of FL* phase in a one-band model is quite challenging because it requires a two-component picture, while all electrons are identical. We adopt the Ancilla Layer Model (ALM) [21] for a mean field description of the FL* phase.
In the ALM framework (as illustrated in Fig. 1 (c)), we introduce two layers of ancilla fermions $\psi_{1}$ and $\psi_{2}$ in addition to the physical electron $c$. $\psi_{1}$ and $\psi_{2}$ is at half filling per spin and they are projected to form a product of rung-singlet state in the end. The physical electron hybridizes with $\psi_{1}$, leading to small Fermi surfaces, while $\psi_{2}$ remains decoupled and forms a spin liquid to be consistent with the Oshikawa-Luttinger theorem [46]. In this picture, $c,\psi_{1}$ form the charge sector while $\psi_{2}$ represent the neutral spinons. In this work, we focus on the charge sector and omit $\psi_{2}$.
@@
$H_{\mathrm{FL*}}=$ $\sum_{\mathbf{k},\sigma}(\epsilon_{c}(\mathbf{k})c^{\dagger}_{\mathbf{k};\sigma}c_{\mathbf{k};\sigma}+\epsilon_{\psi_{1}}(\mathbf{k})\psi^{\dagger}_{1;\mathbf{k};\sigma}\psi_{1;\mathbf{k};\sigma})$ (2)
$+\Phi\sum_{\mathbf{k};\sigma}c_{\mathbf{k};\sigma}^{\dagger}\psi_{1;\mathbf{k};\sigma}+\mathrm{h.c.}~,$
where $\epsilon_{c}(\mathbf{k})$ describes the same large Fermi surface as in Eq. (1), and $\epsilon_{\psi_{1}}(\mathbf{k})$ is the dispersion of the $\psi_{1}$ fermions. When $\Phi$ is finite, the ancilla electrons become part of the Fermi surface and, together with $c$, form small hole-like pockets. In this case, all four Fermi pockets are independent, each with area $A_{\mathrm{FS}}=p/8$, where the factor of 8 arises from the four pockets and spin degeneracy. Notably, the ancilla Fermi Hamiltonian in Eq. (2) naturally explains the Fermi arc behavior. As shown in Fig. 1(b), the inner side of each Fermi pocket is dominated by the physical electron $c$, while the outer side is contributed by the ancilla $\psi_{1}$. The later is not directly visible to spectrum measurement and therefore dark in the ARPES.
@@ III Yamaji effect
In Fig. 1, we illustrate the two different Fermi pockets predicted by the SDW and FL* phases respectively. Under the same bare hopping parameter and a finite $\Phi$, the FL* Fermi pocket is basically the inner half of the SDW pocket, but now the spectrum weight on the back side of the Fermi pocket is negligible because it is dominated by the ancilla $\psi_{1}$. According to (3), the SDW theory always predicts a smaller Yamaji angle than FL* along all the direction, and the distinction becomes most significant along the $\phi=45^{\circ}$ direction. However, it is not clear whether the simple formula in (3) still hold when the spectral weight $Z_{\mathbf{k}}$ has a strong momentum dependence. Therefore, we will firectly calculate the Yamaji angle using a semiclassical method based on our microscopic models in Eq. (1) and Eq. (2).

### v3

@@ Abstract
Recent angle-dependent magnetoresistance measurements in underdoped cuprates have revealed compelling evidence for small hole pockets in the pseudogap regime, including observation of the Yamaji effect in HgBa2CuO4+δ (Chan et al., Nature Physics 21, 1753 (2025)). A key distinction between theories is their predicted Fermi volumes, measured as fractions of the square lattice Brillouin zone: $p/4$ per pocket for spin density wave (SDW) versus $p/8$ for fractionalized Fermi liquid (FL*), where $p$ is the hole doping. We calculate the $c$-axis magnetoresistance $\rho_{zz}(\theta,\phi)$ within the semiclassical Boltzmann formalism for both states, and using the ancilla layer model (ALM) for FL* in a single-band Hamiltonian. The results from the $\text{FL}^{*}$ phase show good consistency with current experimental data. Conversely, the results for the SDW phase are highly sensitive to the ordering momentum along the $z$-direction. An ordering vector of $Q=(\pi,\pi,\pi)$ yields predictions that starkly disagree with the experiment. The only possibility for agreement within the SDW scenario is to assume an ordering momentum of $Q=(\pi,\pi,0)$. However, even in this specific case, the SDW scenario predicts a marginally smaller Yamaji angle at $\phi=0$ than the FL* theory, and a second Yamaji peak near in-plane angle $\phi=45^{\circ}$, which was not observed in the experiment. In reality, the Néel ordering vector is likely uncorrelated between adjacent layers, so that there is no coherent interlayer transport of hole-pocket quasiparticles in the SDW scenario, and consequently no Yamaji effect. Our results support the FL* interpretation of Fermi arcs in the pseudogap phase, and establish Yamaji angle measurements as a discriminatory tool between theoretical models.
@@ I Introduction
The pseudogap phase plays a central role in high-$T_{c}$ cuprates as the parent state from which superconductivity emerges. Despite decades of intensive research, its microscopic origin remains one of the most prominent open questions in the field [54, 39, 24]. A key puzzle concerns the topology of the Fermi surface, which in the pseudogap regime manifests as an open “Fermi arc” in photoemission and scanning tunneling microscopy, rather than a closed, large Fermi surface. Whether these Fermi arcs represent the surviving segments of a large Fermi surface or instead form small Fermi pockets remains highly debated.
In principle, quantum oscillation measurements could discriminate between these scenarios by revealing the Fermi surface geometry directly. However, the very high upper critical fields near optimal doping make such measurements technically challenging. All observed quantum oscillations in the underdoped regime appear in the presence of charge density wave order, which complicates the interpretation of the Fermi surface. Recent progress in high-field transport measurements, particularly $c$-axis angle-dependent magnetoresistance (ADMR) [12, 6], provides a new opportunity to probe the Fermi pockets directly in the pseudogap phase without the presence of field-induced charge density wave order. In particular, Chan et al. studied the compound HgBa2CuO4+δ which has single square lattice CuO2 layers with as AA stacking, making it an ideal platform to study $c$-axis transport, with a simple cosine dispersion along the $c$ axis. In contrast, the earlier study on Nd-LSCO [12] features a staggered stacking of layers, which complicates the features and interpretation of the ADMR [36] (see Appendix F). Chan et al. [6] found that the Yamaji peak positions are consistent with small hole pockets of area $A_{\mathrm{FS}}=p/8$ (measured in units of the Brillouin zone area), rather than a large hole-like Fermi surface of area $A_{\mathrm{Large\,FS}}=(1+p)/2$, where $p$ is the hole doping level. In Ref. [6], the Fermi pocket is approximated as a simple ellipse with area $A_{\mathrm{pocket}}=p/8$, ignoring the Fermi-arc nature of the quasiparticle spectrum. As the spectral weight vanishes on the “back side” of the pocket, it is crucial to revisit whether the Yamaji peak can be reproduced within a more realistic microscopic model of the pseudogap state.
@@
One widely discussed interpretation attributes the Fermi arcs to a spin-density-wave (SDW) order with wavevector near $(\pi,\pi)$ [49, 50, 1, 27, 19, 60, 25]. In this picture, the SDW order coexists with a conventional Fermi liquid (FL) state possessing a large Fermi surface. The SDW order reconstructs the large Fermi surface by folding it with wavevector $(\pi,\pi)$, leading to small hole pockets. This framework provides a satisfactory description in extremely underdoped regime where robust long-range SDW order is present [26]. However, its applicability at higher doping levels is less clear, particularly in the finite-temperature regime where SDW order becomes short-ranged and strongly fluctuating [60, 53, 2].
An alternative scenario does not invoke static SDW order. Instead, it assumes that small Fermi surfaces coexist with a spin-liquid-like background, forming a fractionalized Fermi liquid (FL*) phase [51, 23, 22, 59, 44, 34, 43, 7, 61, 32, 41, 4, 46]. In this case, the small Fermi pocket arises from the binding of holons and spinons into gauge-neutral electron-like quasiparticles, and was predicted to have area $p/8$ [23]. A related binding has been studied in the form of ‘magnetic polarons’ in the low density limit [16, 8], and in the gauge theory context [56].
Another popular approach to the pseudogap is the ‘phase fluctuation’ theory [11, 15, 10, 33, 3, 17, 30, 62, 31, 48, 52, 55, 63, 58]. While this can describe the photoemission Fermi arc spectra in a limited range of temperatures above the superconducting $T_{c}$, it does not have a simple connection to the hole pockets which are the focus of our attention. We also note the holon metal state [28, 45, 9, 5], in which spinless charge $e$ quasiparticles form Fermi pockets of area p/4p/4—these quasiparticles do not have coherent tunneling between layers, and so are inconsistent with ADMR observations.
Thus we focus on the SDW and FL* phases which can be consistent with the observation of the Yamaji effect. The quasiparticles of FL* are gauge neutral, and so can tunnel coherently between layers, as required for the Yamaji effect. For the SDW state, coherent tunneling between layers requires that the SDW ordering correlate between layers. We will consider three dimensional SDW ordering both at $(\pi,\pi,\pi)$ and $(\pi,\pi,0)$, and find that only the $(\pi,\pi,0)$ case could have a Yamaji signal corresponding to the experiments. In neutron scattering experiments on HgBa2CuO4+δ [53, 2], the spin structure factor has negligible dependence on the wavevector in the interlayer direction, indicating negligible spin correlations. This is a serious issue for the SDW explanation of the Yamaji effect, but we will consider it nevertheless.
@@
To check if ADMR can distinguish between the SDW and FL* scenarios, we calculate the $c$-axis magnetoresistance $\rho_{zz}(\theta,\phi)$ within the semiclassical Boltzmann formalism for both cases. A mean-field theory for the FL* state in a single band model, amenable to transport computations, is provided by the ancilla layer model (ALM) [61]. We show that the FL* framework reproduces the observed Yamaji peak as a function of the azimuthal angle $\theta$, at $p=0.1$ in HgBa2CuO4+δ [6]. For in-plane angle $\phi$ close to $0^{\circ}$, the value of the Yamaji $\theta$ is marginally smaller for the SDW theory (assuming $\mathbf{Q}=(\pi,\pi,0))$, but the difference can be magnified by changes in the shape of the FL* hole pocket and the precise definition of the physical electron operator (see Appendix A). Moreover, there is a significant distinction between the two states at $\phi=45^{\circ}$. Here the SDW scenario predicts an additional peak at $\theta\approx 70^{\circ}$, which is absent in both the FL* calculation and in the current experiment [6]. The current experimental result at $\phi=45^{\circ}$ is closer to the FL* scenario than the SDW scenario (even if we assume SDW order with momentum $\mathbf{Q}=(\pi,\pi,0)$). A more definitive conclusion can be reached at larger $\omega_{c}\tau$.
Our results support the FL* interpretation of Fermi arcs in the pseudogap phase, and demonstrate that Yamaji-angle measurements provide a powerful means of distinguishing between competing theoretical scenarios of the pseudogap. We propose future experiments to better distinguish the FL* and SDW scenario based on the data along the in-plane angle $\phi=45^{\circ}$ using higher magnetic fields or cleaner samples.
Figure 1: (a) and (b) shows the Fermi pocket by (a) SDW and (b) FL* by ALM at doping $p=0.1$. In (b), the red color represents the contribution of the physical electron to the spectral weight on the Fermi surface, and purple color represents the contribution from the ancilla fermion $\psi_{1}$. The dark dashed line in (a) indicate the folded Brillouin zone corresponding by SDW $(\pi,\pi)$. (c) An illustration of the ancilla layer model. In the one-band model considered in this work, $c$ and $\psi_{1}$ are the active electrically charged degrees of freedom, while the neutral $\psi_{2}$ fermions decouple and do not contribute to charge transport. (d) Schematic of the polar $\theta$ and azimuthal $\phi$ angles of the magnetic field $\bf B$. And an illustration of the caliper momentum $k_{\mathrm{cal}}$ along the in-plange magnetic field direction $\bf B_{\parallel}$.
@@
SDW metal: In this paper, we mainly consider the $\mathbf{Q}_{2d}=(\pi,\pi)$ SDW order, described by the mean field Hamiltonian:
$H_{\mathrm{SDW}}=$ $\sum_{\mathbf{k},\sigma}(\epsilon_{c}(\mathbf{k}_{2d})-\mu)c^{\dagger}_{\mathbf{k};\sigma}c_{\mathbf{k};\sigma}$
$~~+\sum_{\mathbf{k},\sigma}\Delta\sigma c_{\mathbf{k};\sigma}^{\dagger}c_{\mathbf{k}+\mathbf{Q}_{2d};\sigma}+\mathrm{H.c.}\,.$ (1)
We are using the convention that
$\mathbf{k}=(k_{x},k_{y},k_{z})$ (2)
in a three-dimensional momentum while
$\mathbf{k}_{2d}=(k_{x},k_{y})$ (3)
is its in-plane component, $\epsilon_{c}(\mathbf{k}_{2d})$ denotes the dispersion of free fermions, $\mu$ is the chemical potential, $\mathbf{Q}_{2d}=(\pi,\pi)$ is the SDW ordering wavevector, and $\Delta$ is the SDW order parameter. When $\Delta$ is finite, portions of the original Fermi surface connected by $\mathbf{Q}_{2d}$ are folded together, and a gap opens due to SDW ordering, leading to the formation of small closed pockets. We assume that the SDW order is strong enough to fully gap the electron-like pockets near $(\pi,0)$. The typical Fermi surfaces are shown in Fig. 1(a). In the folded smaller Brillouin zone (BZ), there are two independent Fermi pockets, each with an area of $A=p/4$ in units of the original BZ area, where the factor of 4 in the denominator accounts for the two pockets and two spin degeneracy. Here we have not considered the ordering in the $z$ direction, which will be important in Yamaji effect.
FL* phase: The pseudogap phenomenon is robust across a range of temperatures and doping levels, suggesting that the gap opening and Fermi surface reconstruction may be independent of the symmetry breaking orders. Here we consider another picture that the doped holes (with density $p$) form small pockets on top of the spin moments (with density $1$) from the parent Mott insulator. The exact nature of the spin state for the local moments does not matter much for the hole pockets as long as there is no spin-rotation symmetry breaking. Therefore, we assume the spin moments form a symmetric spin liquid phase. The resulting metallic state is referred fractional Fermi liquid (FL*). The description of FL* phase in a one-band model is quite challenging because it requires a two-component picture, while all electrons are identical. We adopt the Ancilla Layer Model (ALM) [61] for a mean field description of the FL* phase. The reader is referred to recent reviews [4, 46] for further background on FL* and its ALM formulation.
In the ALM framework (as illustrated in Fig. 1 (c)), we introduce two layers of ancilla fermions $\psi_{1}$ and $\psi_{2}$ in addition to the physical electron $c$. The physical electron $c$ hybridizes with $\psi_{1}$, leading to small Fermi surfaces, while $\psi_{2}$ remains decoupled and forms a spin liquid to be consistent with the Oshikawa-Luttinger theorem [40]. In this picture, $c,\psi_{1}$ form the charge sector while $\psi_{2}$ represent the neutral spinons. Both $\psi_{1}$ and $\psi_{2}$ are fixed at half filling per spin, and they are projected to form a product of rung-singlet state in the end to encode the charge and spin degrees of freedom back to the physical band $c$. In this work, we focus on the charge sector and omit $\psi_{2}$.
@@
$H_{\mathrm{FL*}}=$ $\sum_{\mathbf{k},\sigma}(\epsilon_{c}(\mathbf{k}_{2d})-\mu)c^{\dagger}_{\mathbf{k};\sigma}c_{\mathbf{k};\sigma}+\Phi\sum_{\mathbf{k};\sigma}c_{\mathbf{k};\sigma}^{\dagger}\psi_{1;\mathbf{k};\sigma}+\mathrm{H.c.}$ (4)
$+\sum_{\mathbf{k},\sigma}(\epsilon_{\psi_{1}}(\mathbf{k}_{2d})-\mu_{\psi})\psi^{\dagger}_{1;\mathbf{k};\sigma}\psi_{1;\mathbf{k};\sigma}~,$
where $\epsilon_{c}(\mathbf{k}_{2d})$ describes the same large Fermi surface as in Eq. (II), and $\epsilon_{\psi_{1}}(\mathbf{k}_{2d})$ is the dispersion of the $\psi_{1}$ fermions. The chemical potentials $\mu$ and $\mu_{\psi}$ are introduced to fix the average density of $c$ to the physical electron density and that of $\psi_{1}$ to half filling, respectively. When $\Phi$ is finite, the ancilla electrons become part of the Fermi surface and, together with $c$, form small hole-like pockets. In this case, all four Fermi pockets are independent, each with area $A_{\mathrm{FS}}=p/8$, where the factor of 8 arises from the four pockets and spin degeneracy. Notably, the ancilla Fermi Hamiltonian in Eq. (4) naturally explains the Fermi arc behavior. As shown in Fig. 1(b), the inner side of each Fermi pocket is dominated by the physical electron $c$, while the outer side is contributed by the ancilla $\psi_{1}$. The later is not directly visible to spectrum measurement and therefore dark in the ARPES.
@@ III Yamaji effect
In Fig. 1, we illustrate the two different Fermi pockets predicted by the SDW and FL* phases respectively. Under the same bare hopping parameter and a finite $\Phi$, the FL* Fermi pocket is basically the inner half of the SDW pocket, but now the spectrum weight on the back side of the Fermi pocket is negligible because it is dominated by the ancilla $\psi_{1}$. According to (5), the SDW theory always predicts a smaller Yamaji angle than FL* along all the direction, and the distinction becomes most significant along the $\phi=45^{\circ}$ direction. However, it is not clear whether the simple formula in (5) still hold when the spectral weight $Z_{\mathbf{k}}$ has a strong momentum dependence. Therefore, we will firectly calculate the Yamaji angle using a semiclassical method based on our microscopic models in Eq. (II) and Eq. (4).